# MuTox: Universal MUltilingual Audio-based TOXicity Dataset and Zero-shot Detector

### v1

@@ 7 Conclusions
In this paper, we introduced MuTox, a highly multilingual dataset and a massive multilingual toxicity detector that lays the groundwork for the largely unexplored task of multilingual audio-based toxicity detection. Our MuTox dataset enables benchmarking of multilingual audio-based toxicity detection across 21 languages. When compared to cascade tools (speech recognition plus text-based toxicity classifiers), our MuTox classifier demonstrates superior performance on average for all evaluated tasks.
Data [^6] https://github.com/facebookresearch/seamless_communication/tree/main/docs/rai/mutox_README.md and code [^7] https://github.com/facebookresearch/seamless_communication/tree/main/src/seamless_communication/cli/toxicity/mutox/README.md are freely available for further research and development. In future work, we aim to compare MuTox’s performance on the task of added toxicity and plan to extend MuTox to more languages. We also intend to use it to analyze and improve ETOX wordlists, as detailed in the appendix B.. This ongoing work will continue to advance the field of multilingual audio-based toxicity detection.
@@
Annotations were provided by professional annotators, who were informed of the nature of the content to be annotated and were given the opportunity to opt out.
@@
Our dataset may be biased towards text-based toxicity because we pre-selected annotations using this criterion. The use of different text classifiers, as well as a variety of thresholds, mitigates this potential bias.
@@
Our evaluation does not cover unintended biases [^8] https://www.kaggle.com/competitions/jigsaw-unintended-bias-in-toxicity-classification/data, which we intend to cover in future work.
Appendix A Annotations Details
@@
Appendix B Wordlists analysis.
@@
Appendix C Data Card - MuTox
Dataset Summary [^9] We use a template for this data card https://huggingface.co/docs/datasets/v1.12.0/dataset__\__card.html MuTox is a speech and text dataset on 21 languages with toxicity labels. How to use the data You can access links to the data in the README at https://github.com/facebookresearch/seamless_communication/tree/main/docs/rai/mutox_README.md Data Structure Languages: This version of MuTox cover a total of 21 languages: English, Spanish, Arabic, Bengali, Mandarin Chinese, Dutch, French, German, Hindi, Indonesian, Italian, Japanese, Korean, Portuguese, Russian, Swahili, Tagalog, Thai, Turkish, Urdu, Vietnamese Fields available: Speech source, Text source, Toxicity labels Dataset Splits We provide training, dev, devtest and test splits as defined in Table 2. Dataset Creation Preprocessing steps: On speech, we pre-filter audios by length between 2 and 8 seconds for convenience. On text, we pre-selected data based on all criteria described in section 3. Annotations Toxicity labels where produced based on annotation guidelines from section 3. Personal and Sensitive Information: Not applicable Considerations for Using the Data Social Impact of Dataset: The dataset is built to increase toxicity detection quality in speech and text natural language applications. Additional Information Licensing Information We are releasing training, dev and devtest splits under the terms of MIT License Citation Information Marta R. Costa-jussà, Mariano Coria Meglioli, Pierre Andrews, David Dale, Prangthip Hansanti, Elahe Kalbassi, Alex Mourachko, Christophe Ropers, Carleigh Wood, MuTox: Universal MUltilingual Audio-based TOXicity Dataset and Zero-shot Detector, arxiv, 2024
Appendix D Model Card - MuTox
Model Details Person or organization developing model: Developed by FAIR at Meta Model date: November 30, 2023 License: CC-BY-NC 4.0 Intended Use Primary intended uses: Automatic detection of toxicity for speech and text. Information on how to use the model can be found in the Seamless__\__communication repository: https://github.com/facebookresearch/seamless_communication/tree/main/src/seamless_communication/toxicity Primary intended users: Primary users are researchers and speech translation research community. Metrics Model performance measures: AUC, Precision and Recall Data Training and Evaluation data: Detoxy Ghosh et al. (2021), JigSaw Kaggle (2018) and MuTox Corpus, see Table 2 and MuTox data card in appendix B. Ethical Considerations The training data used for model development were mined from various publicly available sources on the web. Although we invested heavily in data cleaning, personally identifiable information may not be entirely eliminated. However, as the model is not generative, it is unlikely that such information could be recovered from the model. We did not specifically study possible biases of the model. Caveats and Recommendations Limitations: Researchers should consider extending data to other languages and more types of toxicity when using the model in a research application. Citation Information Marta R. Costa-jussà, Mariano Coria Meglioli, Pierre Andrews, David Dale, Prangthip Hansanti, Elahe Kalbassi, Alex Mourachko, Christophe Ropers, Carleigh Wood, MuTox: Universal MUltilingual Audio-based TOXicity Dataset and Zero-shot Detector, arxiv, 2024

### v2

@@ 7 Conclusions
In this paper, we introduced MuTox, a highly multilingual dataset and a massive multilingual toxicity detector that lays the groundwork for the largely unexplored task of multilingual audio-based toxicity detection. Our MuTox dataset enables benchmarking of multilingual audio-based toxicity detection across 30 languages. Our MuTox classifier, compared to cascade tools (speech recognition followed by text-based toxicity classifiers) that have similar coverage (more than hundreds of languages), shows superior performance for all evaluated tasks.
Data [^7] Will be revealed after the anonymized review. and code [^8] Will be revealed upon anonymized review. with their corresponding data and model cards are freely available for further research and development. In future work, we primarily aim to explore more complex architectures for training MuTox detector. Additionally, we want to evaluate MuTox’s performance on the task of added toxicity. We also intend to use it to analyze and improve ETOX wordlists, as detailed in the appendix C. This ongoing work will continue to advance the field of multilingual audio-based toxicity detection.
@@
Annotations were provided by professional annotators, who were informed of the nature of the content to be annotated and were given the opportunity to opt out. We understand that the perception of toxicity is subjective, and therefore varies greatly from individual to individual, and from group of individuals to group of individuals. Even within groups of any size, the probability is high that toxicity will be perceived differently from one individual to the next. However, we involved a large pool of annotators with native level in each of the MuTox languages. While our guidelines specify the types of toxicity annotators should look out for, it is ultimately down to annotator judgment to interpret whether a term or phrase is toxic, given the context, and given their personal interpretation.
@@
Our dataset may be biased towards text-based toxicity because we pre-selected annotations using this criterion. The use of different text classifiers, as well as a variety of thresholds, mitigates this potential bias. Moreover, the data is sourced from (e.g. CommonVoice), and it represents a broad and geographically diverse sample.
@@
Our evaluation does not cover unintended biases [^9] https://www.kaggle.com/competitions/jigsaw-unintended-bias-in-toxicity-classification/, which we intend to cover in future work.
Acknowledgements
The authors want to thank Samuel Bell for enriching discussions and his valuable feedback.
Appendix A Languages
arb Modern Standard Arabic Mesopotamian Arabic
ben Bengali Indo-Aryan
bul Bulgarian Balto-Slavic
cat Catalan Romance
ces Czech Balto-Slavic
cmn Mandarin Chinese Sino-Tibetan
dan Danish Germanic
deu German Germanic
ell Greek Hellenic
eng English Germanic
est Estonian Uralic
fas Western Persian Iranian
fin Finnish Uralic
fra French Romance
heb Hebrew Afro-Asiatic
hin Hindi Indo-Aryan
hun Hungarian Uralic
ind Indonesian Austronesian
ita Italian Romance
nld Dutch Germanic
pol Polish Romance
por Portuguese Romance
rus Russian Balto-Slavic
spa Spanish Romance
slk Slovak Balto-Slavic
swh Swahili Atlantic-Congo
tgl Tagalog Austronesian
tur Turkish Turkic
urd Urdu Indo-Aryan
vie Vietnamese Austroasiatic
Table 6: The 30 languages covered in this work.
Appendix B Annotations Details
@@
Appendix C Wordlists analysis.